For two consecutive odd integers, the sum of 3 times the smaller and the larger is 14.

Answer by Maths68(1474)

Let
Smaller odd integer = x+1
Larger odd integer = x+3
Given
Sum of 3 time the smaller and the larger is 14
3(x+1)+x+3=14
3x+3+x+3=14
4x+6=14
4x=14-6
4x=8
4x/4=8/4
x=2


Smaller odd integer = x+1 = 2+1 = 3
Larger odd integer = x+3 = 2+3 = 5